Session: Jun 2005 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Fri Jun 10 21:09:55 2005) Success :

Onsala (Michael Lindqvist at Sat Jun 11 05:09:39 2005) Success :
No known problems.

Torun (K. Borkowski at Mon Jun 13 05:46:56 2005) Success :
No known problems (BBC3 not fixed, as Zsolt communicated, but replaced by higher numbered one).

Effelsberg (A. Kraus at Tue Jun 14 06:34:13 2005) Minor failures : WIND
Heavy rain between UT2015-2050.

Westerbork (Tony Foley at Tue Jun 14 07:04:09 2005) Success :
Wb used in phased array mode. Using 12 telescopes in the tied array (RTC and RTD out for maintenance). No known problems

Jodrell_Bank (Alastair Gunn at Thu Jun 16 10:50:06 2005) Minor failures : RECO
The Lovell Telescope was used. Problem appending to disk pack at start of experiment. Changed pack immediately. First 10 minutes data lost. No other known problems.

Arecibo (TGhosh at Thu Jun 16 12:57:55 2005) Success :
No known problem

Hartebeesthoek (Jonathan Quick at Fri Jun 17 07:28:17 2005) Success :
No known problems